# L-WISE: Boosting human image category learning through model-based image selection and enhancement

### v1

@@ Acknowledgments
In the L-WISE algorithm, we predict the difficulty of each image using its ground truth logit representation ($L_{\text{gt}}$) from a robustified ANN such as ResNet-50. We conducted an additional experiment to compare the ground truth logit with prior state-of-the-art predictors of image difficulty for humans established by Mayo et al. (2023). We apply logistic regression to predict correct v.s. incorrect responses to each image across all participants in our 16-way ImageNet animal classification experiment, using (1) c-score (approximated by the epoch during training at which an image is first correctly predicted (Jiang et al., 2021)), (2) prediction depth (earliest layer upon which a linear probe makes the same prediction as the final output (Baldock et al., 2021)), (3) image-level adversarial robustness (minimum magnitude of image perturbation required to change the network’s prediction), and (4) ground truth logit from both (A) vanilla and (B) robustified ResNet-50 models (see Fig. S4). C-score, prediction depth, and adversarial robustness are implemented following Mayo et al. (2023). The results show that the ground truth logit from robust ResNet-50 ($L_{\text{gt}}$), the metric we use in L-WISE, significantly outperforms all other predictors of image difficulty, including all other metrics combined into one model (“Combined w/o $L_{\text{gt}}$” in Fig. S4). Furthermore, combining all other metrics with $L_{\text{gt}}$ does not improve performance beyond $L_{\text{gt}}$ alone. We also find that using a robustified model greatly improves the predictivity of $L_{\text{gt}}$ and (marginally) adversarial robustness, but not of the c-score or prediction depth.
@@
Figure S4: Robustified ground truth logit is a state-of-the-art predictor of image difficulty for humans, outperforming the c-score, prediction depth, and adversarial epsilon of both vanilla and robustified models. AUC estimates are based on fitted logistic regression models using one or more features listed under each bar. Error bars are 95% confidence intervals for the mean from 10,000 bootstrap replicates. The chance level is AUC=0.5.
@@
Figure S5: Accuracy of image difficulty prediction using ground truth logits from different model types. AUC estimates and 95% confidence interval error bars are generated by the same procedure as in Fig. S4 - however, results are not directly comparable between the two figures as different ANN training runs were used for consistency within each experiment. RN50 = ResNet-50, and $\epsilon$ values in the labels for each bar show the magnitude of the adversarial perturbations during adversarial training. The chance level is AUC=0.5.
@@
Figure S6: Effectiveness of image category enhancement across different guide model types. Each bar shows the mean and 95% confidence interval (by bootstrap) of the rate at which humans choose the original ground truth label, in a 16-way basic animal classification task using ImageNet images. The "Original" bar shows accuracy for unmodified images, and other bars show accuracy of the same participants on images enhanced ($\epsilon=20$) using gradients from the corresponding guide model. RN50 = ResNet-50, and $\epsilon$ values in the labels for each bar show the magnitude of the adversarial perturbations during adversarial training.
@@
Our main results use robustified ResNet-50 as a guide model for perturbations. To evaluate the importance of the choice of guide model, we compared the accuracy of difficulty prediction (Fig. S5) and the effects of enhancement with $eps=20$ (Fig. S6) using 6 different guide models in the 16-way animal classification task. The results show that setting $\epsilon$ to a value of 3 during adversarial training of ResNet-50 yields more accurate difficulty predictions and more effective perturbations than $\epsilon=1$ or $\epsilon=10$ training (consistent with disruption modulation results in Gaziv et al. (2024)), while perturbations guided by a non-adversarially-trained “vanilla” model have negligible effects. Although training ResNet-50 with CutMix improves its robustness to adversarial perturbations (Yun et al., 2019), CutMix-ResNet-50 does not outperform vanilla ResNet-50 in image difficulty prediction and perturbations using it as a guide model do not significantly increase accuracy beyond that on original images. In addition to ResNet-50, we tested the difficulty prediction and image enhancement capabilities of an adversarially trained vision transformer model, XCiT (Ali et al., 2021). Debenedetti et al. (2023) showed that the XCiT architecture is more suitable for adversarial training than the original vision transformer. XCiT generates reasonably accurate image difficulty predictions (on par with the previous state-of-the-art) and generates image perturbations that increase human categorization accuracy by a comparable degree to robustified ResNet-50. For the experiments in Figs. S5 and S6, we used pretrained guide models provided by Gaziv et al. (2024) (Vanilla, $\epsilon=1$, $\epsilon=3$, and $\epsilon=10$ ResNet-50 models), Yun et al. (2019) (CutMix ResNet-50), and Debenedetti et al. (2023) ($\epsilon=4$ XCiT). Examples of images enhanced by each of these guide models with $\epsilon=20$ are displayed in Fig. S7.
@@
Figure S8: Ablation results for image enhancement with ImageNet images. Logit maximization enhancement is effective for images used to train the robustified CNN used as a guide model, and also for held-out validation images (panel A). Logit-max enhancement is more efficient at increasing human accuracy within a given pixel budget ($\epsilon=10$) than enhancement by cross entropy minimization (panel A). The efficacy of logit-max enhancement depends on the difficulty of the original image as estimated by the starting ground truth logit (panel B). In the bar plot of panel B, images were assigned to 4 quadrants based on their ground truth logit values, and for each quadrant the mean difference in accuracy was calculated between original, unmodified images and images enhanced with $\epsilon=10$ (using data from the main 16-way animal classification experiment). The images below each bar illustrate an example image from the category “antelope” drawn from the corresponding difficulty quadrant. All error bars are 95% confidence intervals for the mean from 10,000 bootstrap replicates.
@@
To conduct a limited ablation study on our approach to image enhancement, we conducted an additional 16-way ImageNet animal classification experiment with 20 human participants. This experiment was mostly identical to the 16-way animal classification experiment described in the main text, except there were 6 image conditions instead of 9. Half of the trials used images from the ImageNet validation set (as in the main experiment), and the other half from the training set. Within each training/validation split, one third of the trials were original, unmodified images, one-third were enhanced by maximizing the ground truth logit with $\ell_{2}$ pixel budget $\epsilon=10$, and one-third were enhanced by minimizing the cross-entropy loss with $\epsilon=10$. Results of this experiment are summarized in Fig. S8A. We hypothesized that logit-based enhancement would provide superior results, particularly for images that started off with low cross-entropy loss. We further hypothesized that enhancements would be less effective for training images due to overfitting of the guide model on them. The results show that logit maximization is effective on both training and validation images, and induces a higher increase in accuracy for a given pixel budget $\epsilon$ than cross-entropy minimization. Indeed, cross-entropy minimization significantly increased accuracy only for validation images and not for training images. Unexpectedly, participants were more accurate on original, unmodified training set images than on original, unmodified validation set images. According to Russakovsky et al. (2015), the ImageNet ILSVRC 2012 validation set was collected using the same methodology as the training set, but at a later time. It is therefore plausible that the images and labels in the validation set are drawn from a slightly different distribution than those in the training set, resulting in this accuracy discrepancy.
Figure S9: Plots showing the accuracy trajectory of human participants throughout training/testing in the main dermoscopy learning experiment (panels A-G) and a preceding pilot experiment after which the epsilon tapering schedule was adjusted (panels H-M). All conventions are identical to main-text Fig. 3. There is a statistically significant difference between the test-phase performance of the L-WISE participants and that of the control participants (chi-square test, $p<0.001$) in panel G but not for the pilot experiment in panel M. Notably, the last portion of the training phase does not feature any image enhancements (see Fig. 2F): we suspect that this is the reason for the sudden decline in accuracy in the enhancement group of the pilot experiment (M)
@@
Figure S11: Difficulty prediction and image enhancemement are effective across image domains. Panels A1, B2, and C show the relationship between the ground truth logit from a fine-tuned robustufied ResNet-50 model and the rate at which human participants (from the control group of the ablation study, see Table 1 in the main text) choose the ground truth label during the test phase following a training phase in which they had just attempted to learn the task. Images are binned by ground truth logit to produce the scatter plots, with the number of total trials listed for each bin. Vertical error bars are 95% confidence intervals by bootstrap, and horizontal error bars show the standard deviation. Red lines illustrate fitted logistic regression models. All logistic regression models had statistically significant coefficients for ground truth logit ($p<0.01$). Panels A2 and B2 show the relationship between enhancement $\epsilon$ and the rate at which humans choose the ground truth category. This analysis is limited to the first training trial blocks in the “ET (shuffled)” participant group in the ablation study (main-text Table 1), the only group that viewed enhanced images without monotonically decreasing $\epsilon$. The logistic regression coefficient for $\epsilon$ was statistically significant for dermoscopy images (B2, $p=0.003$) but not for moth photographs (A2, $p=0.14$).
@@
Fig. 3 in the main text shows learning curves (mean accuracy by condition as a function of trial number), and schedules for image difficulty selection and enhancement $\epsilon$, for the moth photograph task: similar plots are shown for the dermoscopy task in Fig. S9 and the histology task in Fig. S10. Panels H-M of Fig. S9 show the results of an early pilot experiment that used image enhancement in isolation (no difficulty selection), in which we suspect the perturbation magnitude $\epsilon$ was set too high causing participants to learn exaggerated features and fail to generalize to natural images with subtler features. This prompted us to switch to the $\epsilon$ schedule we used for our main learning experiments, which starts at $\epsilon=8$ instead of $\epsilon=20$. Panel C of Fig. S10 shows the relationship between the ground truth logit from robust ResNet-50 model and how many of the 7 expert annotators of the MHIST histology dataset Wei et al. (2021) agreed on the same category label: on average, the model is more “confident” in its predictions on images where experts agree to a larger extent.
@@
In addition to the agreement of expert MHIST annotators, the ground truth logit successfully predicts the proportion of human participants who select the correct ground truth label across all tasks we tested. Difficulty prediction results from the 16-way ImageNet task are shown in Fig. 1 in the main text, and from the moth photograph, dermoscopy, and histology tasks in Fig. S11 (Panels A1, B1, and C). For the non-ImageNet tasks, we rely on test-phase data from control group participants who had just learned the tasks in question.
@@
We can also attempt to measure the extent to which images with higher levels of enhancement are easier for novice participants to recognize during the learning tasks (Fig. S11A2,B2). This analysis is limited to the first training trial blocks in the “ET (shuffled)” participant group in the ablation study (main-text Table 1), the only group that viewed enhanced images without monotonically decreasing $\epsilon$. Note that there were 6 discrete $\epsilon$ values (1 per block in the non-shuffled ET condition), and the analysis is complicated by the fact that participants were still learning the task when they made the responses underlying these plots. We are also unable to compare with $\epsilon=0$ unmodified images because participants did not view new unmodified images in the corresponding training blocks. There are no results here for the histology task because the ablation study was conducted only for moth photos and dermoscopy images. In the dermoscopy task, participants respond with the original, correct category label statistically significantly more often when viewing images enhanced with greater $\epsilon$ (Fig. S11B2). A similar trend was not statistically significant for the moth photograph task (Fig. S11B1), perhaps due to the limitations of this particular analysis outlined above. We can observe strong effects of image enhancement in both the moth photograph and dermoscopy tasks by examining the enhancement only (“Enhancement Taper/ET”) arm of the ablation study (Table 1 in the main text): for both tasks, participants in the ET condition had higher test-phase accuracy than participants in the control conditions, and these differences were statistically significant.
@@
To evaluate whether L-WISE has differential effects on human image category learning depending on the image class, we record test-phase precision and recall for each class among L-WISE and control groups in Table S1. The same data are visualized in Fig. 4B. Our experiments are statistically underpowered to detect class-specific differences in performance (as opposed to aggregated performance) - however, we can observe in a coarse sense that the sample means of precision and recall are numerically higher in the L-WISE group across all classes in all tasks. This suggests that overall accuracy improvements attributed to L-WISE are distributed among the various image classes, rather than being the result of isolated improvements in the detection of certain classes.
@@
Nine participants withdrew from the moth photograph category learning experiment. Among them, six had been assigned to the control group, one to the “enhancement taper” group, one to the ”difficulty selection” group, and one to the full L-WISE group. We can calculate the probability of $d=6$ or more participants among the $n=9$ who withdrew being from the control group, under the null hypothesis that the probability of withdrawal is independent of group assignment, using the binomial distribution via Equation 3 below (where $p$ is the probability of being assigned to the control group). Equation 3 evaluates here to a probability of 0.02, indicating that participants who withdrew were significantly more likely to have been assigned to the control group than would be expected if L-WISE assistance had no impact on the probability of withdrawal.
@@
Similarly, in the dermoscopy category learning experiment, 13 participants withdrew, of whom 6 were from the control group. In this case, Equation 3 evaluates to a probability of 0.041, again indicating that participants in the control group withdrew at a significantly higher-than-expected rate. Furthermore, 4 more of the 13 withdrawals were from the “Enhancement Taper (shuffled)” group, which had test-phase accuracy indistinguishable from the control group (see Table 1). None of the withdrawals from the dermoscopy experiment were from the full L-WISE group.
@@
To support our approach to assisting human learners, we demonstrate the ability to enhance category perceptions in response to images using low-norm perturbations. Previous work by Gaziv et al. (2024) showed that an image from one category can be perturbed in a targeted way such that a human perceives it to belong to a different category. Features introduced by these perturbations could be described as “hallucinations:” perceptions (by the model and the human viewer) of objects that are typically not present in the camera’s view. Our image enhancement approach is a special case in the wider realm of categorically targeted image modulation, in which maximization of the ground truth logit perturbs the image such that it becomes a stronger and/or less ambiguous example of its class according to the model’s judgement. Do these perturbations accentuate features that are already present such that they are easier for humans to perceive under challenging conditions, or do they improve human accuracy by hallucinating new features associated with the target category? Subjectively, both phenomena seem to occur: panels C and D in Fig. 1 appear to show bolder contrasts and (in panel C) even the appearance of better camera focus in class-relevant regions of the perturbed images. Panel B in appendix figure Fig. S8 shows a clear example of hallucination, where a semblance of an entire additional “antelope” appears in the foreground of the image. This distinction may be important for education-oriented applications of our enhancement approach, as hallucinations could plausibly impart potentially misleading information to the learner. On the other hand, it is possible that hallucinated features can impart useful and therefore desirable representations of the ground truth class despite departures from a natural image distribution.
@@
We recruited a grand total of 521 participants via the online platform Prolific. All participants live in the United States and are fluent in English (as determined by Prolific). Each participant was eligible to complete each learning experiment only once, to avoid collecting data from participants already familiar with the task.
Our decision regarding the number of participants to recruit for each learning task experimental group (targeting 30 on average) was intended to exceed the requirements of a simple power analysis we conducted following pilot experiments. Pilot experiments showed differences in test-time accuracy between control and either enhancement taper (equivalent to ET in main-text Table 1) or difficulty selection (equivalent to DS in Table 1) participants to be roughly 10%, with a standard deviation of roughly 10% in each group.
@@
Figure S12: Randomized assignment of aliases “Leda,” “Ajax,” “Eris,” and “Tyro” had minimal impact on test-phase accuracy. Each group of four boxplots shows the relative effects of assigning each alias to a specific class from the moth classification experiment. Each individual boxplot indicates the distribution of participant-wise test-phase accuracy z-scores (normalized with mean and standard deviation within each condition separately) among participants with mapping of a specific alias onto a specific class - for example, the left-most boxplot within the right-most group describes the accuracy of participants who saw benign mole images labelled as “Leda.” There is no evidence from one-way ANOVA that the random assignment of aliases to classes influences test-phase performance.

### v2

@@ Acknowledgments
In the L-WISE algorithm, we predict the difficulty of each image using its ground truth logit representation ($L_{\text{gt}}$) from a robustified ANN such as ResNet-50 (see Figs. 1A,  S5, S7, and S14A1,B1,C). We conducted an experiment to compare this ground truth logit score with prior state-of-the-art predictors of image difficulty for humans established by Mayo et al. (2023). We apply logistic regression to predict correct v.s. incorrect responses to each (original, unmodified) image across all participants in our 16-way ImageNet animal classification experiment, using (1) c-score (approximated by the epoch during training at which an image is first correctly predicted (Jiang et al., 2021)), (2) prediction depth (earliest layer upon which a linear probe makes the same prediction as the final output (Baldock et al., 2021)), (3) image-level adversarial robustness (minimum magnitude of image perturbation required to change the network’s prediction), and (4) ground truth logit from both (A) vanilla and (B) robustified ResNet-50 models (see Fig. S6). C-score, prediction depth, and adversarial robustness are implemented following Mayo et al. (2023). The results show that the ground truth logit from robust ResNet-50 ($L_{\text{gt}}$), the metric we use in L-WISE, significantly outperforms all other predictors of image difficulty, including all other metrics combined into one model (“Combined w/o $L_{\text{gt}}$” in Fig. S6). Furthermore, combining all other metrics with $L_{\text{gt}}$ does not improve performance beyond $L_{\text{gt}}$ alone. We also find that using a robustified model rather than a “vanilla” model to generate each metric greatly improves the predictivity of $L_{\text{gt}}$ and (marginally) adversarial robustness, but not of the c-score or prediction depth.
@@
Figure S6: Robustified ground truth logit is a state-of-the-art predictor of image difficulty for humans, outperforming the c-score, prediction depth, and adversarial epsilon of both vanilla and robustified models. AUC estimates are based on fitted logistic regression models using one or more features listed under each bar, with stratified 500-fold cross-validation. Error bars are 95% confidence intervals for the mean from 10,000 bootstrap replicates. The chance level is AUC=0.5.
@@
Figure S7: Accuracy of image difficulty prediction using ground truth logits from different model types. AUC estimates and 95% confidence interval error bars are generated by the same procedure as in Fig. S6 - however, results are not directly comparable between the two figures as different ANN training runs were used for consistency within each experiment. RN50 = ResNet-50, and $\epsilon$ values in the labels for each bar show the magnitude of the adversarial perturbations during adversarial training.
[…]
@@
Figure S9: Effectiveness of image category enhancement across different guide model types. Each bar shows the mean and 95% confidence interval (by bootstrap) of the rate at which humans choose the original ground truth label, in a 16-way basic animal classification task using ImageNet images. The “Original” bar shows accuracy for unmodified images, and other bars show accuracy of the same participants on images enhanced ($\epsilon=20$) using gradients from the corresponding guide model. RN50 = ResNet-50, and $\epsilon$ values in the labels for each bar show the magnitude of the adversarial perturbations during adversarial training.
@@
Our main results use robustified ResNet-50 as a guide model for generating image perturbations. To evaluate the importance of the choice of guide model, we compared the accuracy of difficulty prediction (Fig. S7) and the effects of enhancement with $\epsilon=20$ (Fig. S9) using 6 different guide models in the 16-way animal classification task. The results show that setting $\epsilon$ to a value of 3 during adversarial training of ResNet-50 yields more accurate difficulty predictions and more effective perturbations than $\epsilon=1$ or $\epsilon=10$ training (consistent with disruption modulation results in Gaziv et al. (2023)), while perturbations guided by a non-adversarially-trained “vanilla” model have negligible effects. Model accuracy seems to be less important than robustness, at least for difficulty prediction: ground truth logits from ResNet-50 models at early epochs of adversarial training on ImageNet, which have lower image classification accuracy than models at later epochs, can still yield accurate image difficulty predictions (Fig. S8A). Models at early training epochs may also be able to generate high-quality image enhancement perturbations (Fig. S8B), although we did not study this in our experiments with human participants. Although training ResNet-50 with CutMix improves its robustness to adversarial perturbations (Yun et al., 2019), CutMix-ResNet-50 does not outperform vanilla ResNet-50 in image difficulty prediction and perturbations using it as a guide model do not significantly increase accuracy beyond that on original images. In addition to ResNet-50, we tested the difficulty prediction and image enhancement capabilities of an adversarially trained vision transformer model, the Cross-Covariance Image Transformer (XCiT) (Ali et al., 2021). Debenedetti et al. (2023) showed that the XCiT architecture is more suitable for adversarial training than the original vision transformer. XCiT generates reasonably accurate image difficulty predictions (on par with the previous state-of-the-art) and generates image perturbations that increase human categorization accuracy by a comparable degree to robustified ResNet-50. For the experiments in Figs. S7 and S9, we used pretrained guide models provided by Gaziv et al. (2023) (Vanilla, $\epsilon=1$, $\epsilon=3$, and $\epsilon=10$ ResNet-50 models), Yun et al. (2019) (CutMix ResNet-50), and Debenedetti et al. (2023) ($\epsilon=4$ XCiT). Examples of images enhanced by each of these guide models with $\epsilon=20$ are displayed in Fig. S10.
@@
Figure S11: Ablation results for image enhancement with ImageNet images. Logit maximization enhancement is effective for images used to train the robustified CNN used as a guide model, and also for held-out validation images (panel A). Logit-max enhancement is more efficient at increasing human accuracy within a given pixel budget ($\epsilon=10$) than enhancement by cross-entropy minimization (panel A). The efficacy of logit-max enhancement depends on the difficulty of the original image as estimated by the starting ground truth logit (panel B). In the bar plot of panel B, images were assigned to 4 quadrants based on their ground truth logit values, and for each quadrant the mean difference in accuracy was calculated between original, unmodified images and images enhanced with $\epsilon=10$ (using data from the main 16-way animal classification experiment). The images below each bar illustrate an example image from the category “antelope” drawn from the corresponding difficulty quadrant. All error bars are 95% confidence intervals for the mean from 10,000 bootstrap replicates.
@@
As a limited ablation study on our approach to image enhancement, we conducted an additional 16-way ImageNet animal classification experiment with 20 human participants. This experiment was mostly identical to the 16-way animal classification experiment described in the main text, except there were 6 image conditions instead of 9. Half of the trials used images from the ImageNet validation set (as in the main experiment), and the other half from the training set. Within each training/validation split, one third of the trials were original, unmodified images, one-third were enhanced by maximizing the ground truth logit with $\ell_{2}$ pixel budget $\epsilon=10$, and one-third were enhanced by minimizing the cross-entropy loss with $\epsilon=10$. The results of this experiment are summarized in Fig. S11A. We hypothesized that logit-based enhancement would provide superior results, particularly for images that started off with low cross-entropy loss. We further hypothesized that enhancements would be less effective for training images due to overfitting of the guide model on them. The results show that logit maximization is effective on both training and validation images, and induces a larger increase in accuracy for a given pixel budget $\epsilon$ than cross-entropy minimization. Indeed, cross-entropy minimization significantly increased accuracy only for validation images and not for training images. Unexpectedly, participants were more accurate on original, unmodified training set images than on original, unmodified validation set images. According to Russakovsky et al. (2015), the ImageNet ILSVRC 2012 validation set was collected using the same methodology as the training set, but at a later time. It is therefore plausible that the images and labels in the validation set are drawn from a slightly different distribution than those in the training set, resulting in this accuracy discrepancy.
Figure S12: Plots showing the accuracy trajectory of human participants throughout training/testing in the main dermoscopy learning experiment (panels A-G) and a preceding pilot experiment after which the epsilon tapering schedule was adjusted (panels H-M). All conventions are identical to main-text Fig. 3. There is a statistically significant difference between the test-phase performance of the L-WISE participants and that of the control participants ($\chi^{2}(1)$ test, $p<0.001$) in panel G but not for the pilot experiment in panel M. Notably, the last portion of the training phase does not feature any image enhancements (see Fig. 2F): we suspect that this is the reason for the sudden decline in accuracy in the enhancement group of the pilot experiment (M).
@@
Figure S14: Difficulty prediction and image enhancemement are effective across image domains. Panels A1, B1, and C show the relationship between the ground truth logit from a fine-tuned robustufied ResNet-50 model and the rate at which human participants (from the control groups) choose the ground truth label during the test phase following a training phase in which they had just attempted to learn the task. Images are binned by ground truth logit to produce the scatter plots, with the number of total trials listed for each bin. Vertical error bars are 95% confidence intervals by bootstrap, and horizontal error bars show the standard deviation within each bin. The curved lines illustrate fitted logistic regression models. All logistic regression models had statistically significant coefficients for ground truth logit ($p<0.001$ from Wald statistic). Panels A2 and B2 show the relationship between enhancement $\epsilon$ and the rate at which humans choose the ground truth category. This analysis is limited to the first training trial blocks in the “ET (shuffled)” participant group in the ablation study (main-text Table 1), the only group that viewed enhanced images without $\epsilon$ monotonically decreasing over time. The logistic regression coefficient for $\epsilon$ was statistically significant ($p<0.05$ from Wald statistic) for both moth photographs (A2) and dermoscopy images (B2).
@@
Fig. 3 in the main text shows learning curves (mean smoothed accuracy by condition as a function of trial number), and schedules for image difficulty selection and enhancement $\epsilon$, for the moth photograph task: similar plots are shown for the dermoscopy task in Fig. S12 and the histology task in Fig. S13. Panels H-M of Fig. S12 show the results of an early pilot experiment that used image enhancement in isolation (no difficulty selection), in which we suspect the perturbation magnitude $\epsilon$ was set too high causing participants to learn exaggerated features and fail to generalize to natural images with subtler features. This prompted us to switch to the $\epsilon$ schedule we used for our main learning experiments, which starts at $\epsilon=8$ instead of $\epsilon=20$. Panel C of Fig. S13 shows the relationship between the ground truth logit from robust ResNet-50 model and how many of the 7 expert annotators of the MHIST histology dataset Wei et al. (2021) agreed on the same category label. On average, the model is more “confident” in its predictions (higher ground truth logit) on images where experts are more in agreement with each other.
@@
In addition to the agreement of expert MHIST annotators, the ground truth logit successfully predicts the proportion of human participants who select the correct ground truth label across all tasks we tested. Difficulty prediction results from the 16-way ImageNet task are shown in Fig. 1 in the main text, and from the moth photograph, dermoscopy, and histology tasks in Fig. S14 (Panels A1, B1, and C). For this analysis in the non-ImageNet tasks, we rely on test-phase data from control group participants who had just learned the tasks in question.
@@
We can also attempt to measure the extent to which images with higher levels of enhancement are easier for novice participants to recognize during the learning tasks (Fig. S14A2,B2). This analysis is limited to the first training trial blocks in the “ET (shuffled)” participant group in the ablation study (main-text Table 1), the only group that viewed enhanced images without $\epsilon$ monotonically decreasing over time. Note that there were 6 discrete $\epsilon$ values (1 per block in the non-shuffled ET condition), and the analysis is complicated by the fact that participants were still learning the task when they made the responses underlying these plots. We are also unable to compare with $\epsilon=0$ unmodified images because participants did not view new unmodified images in the corresponding training blocks. There are no results here for the histology task because the ablation study was conducted only for moth photos and dermoscopy images. In both the moth task and the dermoscopy task, participants respond with the original, correct category label statistically significantly more often when viewing images enhanced with greater $\epsilon$ (Fig. S14B1,B2).
@@
To evaluate whether L-WISE has differential effects on human image category learning depending on the image class, we record test-phase precision and recall for each class among L-WISE and control groups in Table S1. The same data are visualized in Fig. 4B. Our experiments are statistically underpowered to detect class-specific differences in performance (as opposed to aggregated performance) - however, we can observe in a coarse sense that the sample means of precision and recall are numerically higher in the L-WISE group across all classes in all tasks. This suggests that overall accuracy improvements attributed to L-WISE are distributed among the various image classes, rather than being the result of isolated improvements in the detection of a subset of classes.
@@
Nine participants withdrew from the moth photograph category learning experiment. Among them, six had been assigned to the control group, one to the “enhancement taper” group, one to the ”difficulty selection” group, and one to the full L-WISE group. We can calculate the probability of $d=6$ or more participants among the $n=9$ who withdrew being from the control group, under the null hypothesis that the probability of withdrawal is independent of group assignment, using the binomial distribution via Equation 3 below (where $p$ is the probability of being assigned to the control group). Equation 3 evaluates here to a probability of $p=0.02$, indicating that participants who withdrew were significantly more likely to have been assigned to the control group than would be expected if L-WISE assistance had no impact on the probability of withdrawal.
@@
Similarly, in the dermoscopy category learning experiment, 13 participants withdrew, of whom 6 were from the control group. In this case, Equation 3 evaluates to a probability of $p=0.041$, again indicating that participants in the control group withdrew at a significantly higher-than-expected rate. Furthermore, 4 more of the 13 withdrawals were from the “Enhancement Taper (shuffled)” group, which had test-phase accuracy indistinguishable from the control group (see Table 1). None of the withdrawals from the dermoscopy experiment were from the full L-WISE group.
@@
To support our approach to assisting human learners, we demonstrate the ability to enhance category percepts in images using low-norm perturbations. Previous work by Gaziv et al. (2023) showed that an image from one category can be perturbed in a targeted way such that a human perceives it to belong to a different category. Features introduced by these disruptive perturbations could be described as “hallucinations:” perceptions (by the model and the human viewer) of objects that are not present in the camera’s view. Our image enhancement approach is a special case in the wider realm of categorically targeted image modulation, in which maximization of the ground truth logit perturbs the image such that it becomes a stronger and/or less ambiguous example of its class according to the model’s judgement. Do these perturbations accentuate features that are already present such that they are easier for humans to perceive under challenging conditions, or do they improve human accuracy by hallucinating new features associated with the target category? Subjectively, both phenomena seem to occur: panel B2 in Fig. 1 appears to show bolder contrasts and exaggerated features in class-relevant regions of the perturbed images. Panel B (image farthest to the left) in Appendix Fig. S11, however, shows a clear example of hallucination, where a semblance of an entire additional “antelope” appears in the foreground of the image. This distinction may be important for education-oriented applications of our enhancement approach, as hallucinations could plausibly impart potentially misleading information to the learner. On the other hand, it is possible that hallucinated features can impart useful and therefore desirable representations of the ground truth class despite departures from a natural image distribution.
@@
We recruited a grand total of 521 participants via the online platform Prolific. All participants lived in the United States and were fluent in English (as determined by Prolific). Each participant was eligible to complete each learning experiment only once, to avoid collecting data from participants already familiar with a given task.
Our decision regarding the number of participants to recruit for each learning task experimental group (targeting 30 on average) was intended to exceed the requirements of a simple power analysis we conducted following pilot experiments. Pilot experiments showed differences in test-time accuracy between control and either enhancement taper (equivalent to ET in main-text Table 1) or difficulty selection (equivalent to DS in Table 1) participants to be roughly 10%, with a standard deviation in accuracy of roughly 10% in each group.
@@
Figure S15: Randomized assignment of aliases “Leda,” “Ajax,” “Eris,” and “Tyro” had minimal impact on test-phase accuracy in the moth species classification task. Each group of four boxplots shows the relative effects of assigning each alias to a specific class from the moth classification experiment. Each individual boxplot indicates the distribution of participant-wise test-phase accuracy z-scores (normalized with mean and standard deviation within each condition separately) among participants with mapping of a specific alias onto a specific class - for example, the left-most boxplot within the right-most group describes the accuracy of participants who saw images of the moth species Idaea tacturata labelled as “Eris.” There is no evidence from one-way ANOVA that the random assignment of aliases to classes influences test-phase performance.